OpenStudy (anonymous):

A boat traveled downstream a distance of 80 mi and then came right back. If the speed of the current was 20 mph and the total trip took 5 hours and 20 minutes, find the average speed of the boat relative to the water.

OpenStudy (radar):

Use the d=rt formula d = distance in miles = 80 (given) r = rate in mph (to be solved t= time in hours for round trip = 5 1/3 Hours (given

OpenStudy (radar):

t = d/r For the trip down stream t=80/(r +20) down stream time t=80/(r-20) up stream time the total travel time is 5 1/3 hour so we know 80/(r+20) + 80/(r-20) = 5 1/3 or 16/3 Now solve for r.

OpenStudy (radar):

add the two fractions whose LCD is (r-20)(r+20) You will come up with a quadratic equation that can be factored and solved